OpenStudy (anonymous):

a close cylindrical tank has a capacity of 576.56 cu. m. what must the minimum surface area of the tank? a. 412. 60 cu. m. b 383. 40 cu. m. c 516. 32 cu. m. d 218. 60 cu. m.

OpenStudy (agent0smith):

First you need formulae for the volume and surface area of a cylinder: http://math.about.com/od/formulas/ss/surfaceareavol_3.htm First, why are the answers all in cubic metres? Surface area should be in metres squared.

OpenStudy (anonymous):

i dont know y im nose bleed in english hahaha

OpenStudy (anonymous):

v=pi*r^2 h diff it get dh/dr then diff the surface area replace dr to dh using equation from vol and then equate with 0 get r/h get answer

OpenStudy (anonymous):

any problem in the given theory

OpenStudy (agent0smith):

Ah well, i'm guessing they're meant to be sq cm. From the volume formula, we can get either r in terms of h, or h in terms of r. h in terms of r is probably easier, given the SA equation. \[V= 576.56 = \pi r ^{2}h \] \[h = \frac{ V }{ \pi r ^{2} }\] Now put that into the surface area equation and simplify a bit: \[SA = 2\pi r ^{2}+\frac{ 2V }{ r }\] Then differentiate and set SA' = 0.

OpenStudy (anonymous):

ano po yun answer?

OpenStudy (agent0smith):

\[SA' = 0 = 4\pi r - \frac{ 2V }{ r ^{2}}\] rearrange for r: \[r = \sqrt[3]{\frac{ V }{ 2\pi }} \] so r = 4.51 and now sub that into your SA equation.